The radius of the aorta is about 1.1 cm , and the blood passing through it has a speed of about 48 cm/s . Coefficient of viscosity for the whole blood (37∘) is η = 4×10−3Pa⋅s.Calculate the pressure drop per cm along the aorta.
Answer #1

Pressure drop is given by

Q = \frac{\delta P \pi r^4}{8\eta l}

\pi r^2 v = \frac{\delta P \pi r^4}{8\eta l}

Plug in all values we will have

P = 1.27*10^5 Pa